OpenStudy (anonymous):

what is the x-intercept of the parabola. y=-3x^2+12x-9

OpenStudy (anonymous):

is it (9,3)

OpenStudy (anonymous):

put y=0 and solve for x

OpenStudy (anonymous):

The x-intercepts are the points where the coordinate of the parabola equals 0, i.e. y = 0. So, equating y to 0, we get: \[-3x^2 + 12x - 9 = 0\] \[-3x^2 +9x +3x - 9 = 0\] \[(x-3)(3x-3)=0\] \[x=3, x=1\] Those are the x-intercepts.